OpenStudy (babyslapmafro):

Please help me find an exact, simplified answer to the following... 11(-pi/4)-22sin(-pi/4)

OpenStudy (babyslapmafro):

\[11\frac{ -\pi }{ 4 }-22\sin(\frac{ -\pi }{ 4 })\]